Montserrat Volcano Observatory

Daily Report
Report for the period 5 p.m. 27 July
to 5 p.m. 28 July 1998

Activity at the volcano was lower today than yesterday. 4 volcano-tectonic events and 8 rockfall signals triggered the network.

COSPEC measurements were made in conjunction with another method for measuring gas output from the volcano. This alternative method uses infrared light from the sun to measure the amount of various different gases in the volcanic plume. Results from these experiments should be available in a later report.

GPS occupations have taken place today at Blakes, Harris, Hermitage and at Old Towne. The data are currently being processed.

Dust levels were low at all stations except the monitoring site in Salem.

Residents should continue listening to Radio Montserrat in case there are sudden developments in the state of the volcano.

Montserrat Volcano Observatory